OpenStudy (anonymous):

lim 5^t - 3^t / t (l'hospital's rule section) t->0

myininaya (myininaya):

Is that \[\lim_{t \rightarrow 0}(5^t-\frac{3^t}{t}) \text{ or } \lim_{t \rightarrow 0}\frac{5^t-3^t}{t}\]

myininaya (myininaya):

I think you might mean the second one in which case plugin' in 0 will give you 0/0 which means we can use l'hospital's rule

myininaya (myininaya):

\[\lim_{t \rightarrow 0}\frac{5^t-3^t}{t}=\lim_{t \rightarrow 0}\frac{(5^t-3^t)'}{(t)'}\]

myininaya (myininaya):

I'm assuming your trouble is with the top?

myininaya (myininaya):

\[\text{ if } y=a^x \text{ and } a>0 \text{ then } y'=a^x \cdot \ln(a)\]

OpenStudy (anonymous):

yes, so i have \[\frac{5^{t}\ln 5 - 3^{t} \ln 3 }{1}\]

myininaya (myininaya):

ok great now plug in 0

OpenStudy (anonymous):

i get ln 5 - ln 3 which is about .51 on calc... but answer given is \[\ln \frac{5}{3}\]

myininaya (myininaya):

\[\ln(5)-\ln(3)=\ln(\frac{5}{3})\]

myininaya (myininaya):

by properties of log

OpenStudy (anonymous):

oh yeah.

OpenStudy (anonymous):

ty

myininaya (myininaya):

np